OpenStudy (anonymous):

FAN &.MEDAL A toy storechanged the price of a puzzle from $5.00 to $8.00. What was the percent of increase of the new price 60% 38% 160% 63%

OpenStudy (anonymous):

The price increase by $3 compared to the original of $5

OpenStudy (anonymous):

compared to $5, how much s $3?

OpenStudy (acxbox22):

8-5/5=3/5=0.6=60%

OpenStudy (mokeira):

increase= (8.00=5.00)=$3.00 percentage increase \[\frac{ 3 }{5 }\times100=simplify\]

OpenStudy (acxbox22):

it is difference between original and new 8-5 ------------------------------- ==== ----------- original price 5

OpenStudy (anonymous):

Thnx